OpenStudy (anonymous):

Give a unique example of each: 1. Part 1: An absolute value equation with two solutions 2. Part 2: An absolute value equation with one solution 3. Part 3: An absolute value equation with no solutions

OpenStudy (anonymous):

please helppp!!

OpenStudy (anonymous):

1) |x+2| = 3 2) |x+2| = 0 3) |x+2| = -3